### v2

# When Respondents Don’t Care Anymore: Identifying the Onset of Careless Responding

Abstract
Questionnaires in the behavioral and organizational sciences tend to be lengthy: survey measures comprising hundreds of items are the norm rather than the exception. However, literature suggests that the longer a questionnaire takes, the higher the probability that participants lose interest and start responding carelessly. Consequently, in long surveys a large number of participants may engage in careless responding, posing a major threat to internal validity. We propose a novel method for identifying the onset of careless responding (or an absence thereof) for each participant. It is based on combined measurements of multiple dimensions in which carelessness may manifest, such as inconsistency and invariability. Since a structural break in either dimension is potentially indicative of carelessness, the proposed method searches for evidence for changepoints along the combined measurements. It is highly flexible, based on machine learning, and provides statistical guarantees on its performance. An empirical application on data from a seminal study on the incidence of careless responding reveals that the reported incidence has likely been substantially underestimated due to the presence of respondents that were careless for only parts of the questionnaire. In simulation experiments, we find that the proposed method achieves high reliability in correctly identifying carelessness onset, discriminates well between careless and attentive respondents, and captures a variety of careless response types, even when a large number of careless respondents are present. Furthermore, we provide freely available open source software to enhance accessibility and facilitate adoption by empirical researchers.
Keywords: careless responding, changepoint detection, machine learning, survey methodology
Introduction
Research in the behavioral and organizational sciences often involves the administration of lengthy (online) self-report questionnaires. For instance, common personality measures tend to consist of hundreds of items, such as the Revised NEO Personality Inventory (Costa & McCrae,, 1992) with 240 items or the Minnesota Multiphasic Personality Inventory-2 Restructured Form (Ben-Porath & Tellegen,, 2008) with 338 items. Even if one does not use such extensive measures, the number of items can easily reach three digits by including several shorter measures. However, it is well known that questionnaire length can have a concerning adverse effect on measurement accuracy: Respondents may experience fatigue or boredom as they progress through a lengthy questionnaire, which can provoke careless responding (e.g., Bowling et al.,, 2021; Gibson & Bowling,, 2020; Galesic & Bosnjak,, 2009), in particular in online administrations (e.g., Johnson,, 2005; Meade & Craig,, 2012). We call such participants partially careless respondents as they initially provide accurate responses, but resort to careless responding after a certain item and remain careless for the remainder of the questionnaire.
According to the definition of Ward & Meade, (2023), “careless responding occurs when participants are not basing their response on the item content, and it can occur when a respondent does not read an item, does not understand an item, or is unmotivated to think about what the item is asking”.11endnote: 1Perhaps more commonly used, Huang et al., (2012) define careless responding as “a response set in which the respondent answers a survey measure with low or little motivation to comply with survey instructions, correctly interpret item content, and provide accurate responses”. We prefer the definition of Ward & Meade, (2023), as it highlights that careless responding may not be intentional (cf. Huang et al.,, 2012; Ward & Pond,, 2015; Ward et al.,, 2017). Also note that Huang et al., (2012) refer to careless responding as insufficient effort responding in their definition. Other synonyms for careless responding are participant inattention (Maniaci & Rogge,, 2014), inconsistent responding (Greene,, 1978), protocol invalidity (Johnson,, 2005), and random responding (e.g., Beach,, 1989; Berry et al.,, 1992; Emons,, 2008). Notably, Schroeders et al., (2022) point out that the latter—random responding—may be a misnomer since careless responding can also arise in non-random patterns (e.g., a recurring sequence of 1-2-3-4-5). Careless responding has been identified as a major threat to the validity of research findings through a variety of psychometric and statistical issues such as lower scale reliability, errors in statistical hypothesis testing, and deteriorated model fit (e.g., Arias et al.,, 2020; Huang et al.,, 2012, 2015; Kam & Meyer,, 2015). It is widely prevalent (e.g., Bowling et al.,, 2016; DeSimone et al.,, 2015; Curran,, 2016; Meade & Craig,, 2012), suspected to be present in all questionnaire data (Ward & Meade,, 2023), and already incidence rates as low as 5% can be problematic (Arias et al.,, 2020; Credé,, 2010). Often viewed as merely an insidious source of measurement error, careless responding can be a valuable source of information about study participants and the study itself. For instance, carelessness has been linked to participant personality (e.g., Bowling et al.,, 2016; DeSimone et al.,, 2020) and design factors such as study length (e.g., Bowling et al.,, 2021) and item wording. In particular negatively-worded items are prone to misresponses (e.g., Swain et al.,, 2008; Weijters & Baumgartner,, 2012) triggered by carelessness. Consequently, it is widely recommended to carefully screen survey data for the presence of careless responding (e.g., DeSimone et al.,, 2015; Huang et al.,, 2015).
Numerous methods for identifying careless respondents after questionnaire administration (so-called screeners) have been proposed, for instance consistency indicators such as psychometric synonyms (Meade & Craig,, 2012), longstring indices (Johnson,, 2005), or multivariate outlier analyses (e.g., Curran,, 2016). We refer to Arthur et al., (2021), DeSimone et al., (2015), and Ward & Meade, (2023) for detailed reviews.22endnote: 2A popular pre-administration method for the detection of careless responding is the inclusion of detection items (e.g., Table 1 in Arthur et al.,, 2021). Such detection items are based on the presumption that an attentive respondent will respond in a specific way, while careless respondent may fail to do so. For instance, it is expected that an attentive respondent would strongly disagree to so-called bogus items such as “I am paid biweekly by leprechauns” (Meade & Craig,, 2012). A careless respondent may accidentally “agree” to this item as a consequence of inattention. Alternative types of detection items are instructed items and self-report items (Arthur et al.,, 2021). We discuss detection items and preventive measures against carelessness in detail in a later section. More recently, Schroeders et al., (2022), Alfons & Welz, (2024), and Ozaki, (2024) have studied machine learning techniques for identifying careless respondents. Still, commonly used screeners tend to suffer from three major limitations, which we discuss in the following.
First, extant methods commonly follow the implicit assumption that careless respondents respond carelessly to all questionnaire items or at least a vast majority of items. In particular, they indicate whether a respondent engaged in careless responding, but they typically do not provide any indication regarding which items are affected. Although there exist many comparative studies (e.g., Curran,, 2016; DeSimone & Harms,, 2018; Goldammer et al.,, 2020; Huang et al.,, 2012) on such screeners, their behavior under partial careless responding is not well-understood. This is at odds with the substantial evidence that many careless respondents tend to begin a survey as attentive and truthful respondents, but may resort to careless responding due to fatigue or boredom as a survey progresses (Bowling et al.,, 2021; Brower,, 2018; Galesic & Bosnjak,, 2009; Gibson & Bowling,, 2020; Ward et al.,, 2017). In general, Ward & Meade, (2023) believe it is rare that careless respondents respond carelessly from the beginning to the end of a questionnaire.
Second, common detection methods are typically designed to detect a single specific type of carelessness. For instance, the longstring index of Johnson, (2005) counts the maximum number of consecutive identical responses, which is intended to capture straightlining behavior. However, careless responding may manifest in multiple distinctive ways (e.g., Curran,, 2016; DeSimone et al.,, 2015; DeSimone & Harms,, 2018; Edwards,, 2019; Ward & Meade,, 2023), implying that a screener that is designed to detect one type of carelessness may fail to detect other types. It is therefore recommended to combine multiple screeners to detect various types of careless respondents (e.g., Goldammer et al.,, 2020; Huang et al.,, 2012; Meade & Craig,, 2012; Ward & Meade,, 2023).
Third, cutoff values for flagging careless respondents are often ad-hoc and lack statistical foundation. Screeners typically calculate a score for each participant, with a participant being flagged as careless if their score is sufficiently extreme (i.e., beyond a certain cutoff value). Ideally, such a cutoff value results in high sensitivity (flagging true careless respondents as careless) and high specificity (not flagging true attentive respondents as careless). For certain screeners, various cutoff values have been suggested in the literature, but individual suggestions vary substantially. For instance, for the same screeners, Goldammer et al., (2020) argues for different cutoff values than Curran, (2016) or Johnson, (2005), which suggests that the results for specific cutoff values may depend on factors such as the population or the constructs being measured (cf. Meade & Craig,, 2012, p. 452). For a specific screener, it would be desirable to obtain a cutoff value with statistical guarantees on sensitivity and specificity that does not depend on population or study characteristics.
In wake of these three issues with extant methods, we propose a novel screener for the identification of careless responding that is designed to overcome these issues. We achieve this by focusing on a more nuanced question. Motivated by the prominence of partial careless responding due to survey fatigue or boredom, we focus on identifying at which point in the survey a given participant becomes careless (if at all). Specifically, our method estimates the item at which carelessness onsets in each questionnaire participant, or an absence thereof. Hence, we call our method careless onset detection in extensive rating-scale surveys (CODERS). To capture various types of carelessness, CODERS can combine multiple dimensions of evidence for careless responding to construct a score that, for each item, measures if a given respondent has started responding carelessly by that item. It achieves this by searching for changepoints along a multidimensional series of indices that are potentially indicative of carelessness, such as measures of response consistency or variability.
We argue that the notion of a changepoint is intuitive when studying partial carelessness: Once a participant starts responding carelessly, we expect a structural break in their responses. In particular, such a respondent may abandon content-based responding and resort to careless response types, while no such break occurs in the absence of carelessness. CODERS is highly flexible as it does not assume a statistical model, nor does it require predefining what types of careless response types may exist, and it is primarily intended for lengthy multi-scale surveys. Moreover, CODERS builds on nonparametric estimation theory for cutoff values and statistical guarantees regarding sensitivity and specificity.
Re-examining data from a a seminal study on careless responding in Big 5 measurements (Johnson,, 2005), we demonstrate novel insights that can be obtained by estimating onset items of partially careless responding. Most notably, we find strong evidence that Johnson, (2005) considerably underestimates the incidence of careless respondents due to hitherto undetected presence of partially careless respondents. Through extensive simulation experiments, we find that the proposed method CODERS accurately estimates carelessness onset (or a lack thereof) and thereby achieves excellent sensitivity and specificity in detecting careless responding.
To the best of our knowledge, CODERS is the first attempt to systematically detect the onset of careless responding. Assuming that participants respond accurately in the beginning of the survey, CODERS is designed to segment the responses of each respondent into a segment of accurate responses and—if a changepoint was identified—a segment of careless responses. With this knowledge, researchers can restrict their primary analyses to the segments of accurate responses (by treating the careless responses as missing values) without having to discard all responses of a partially careless respondent. In addition, researchers can separately study the segments of careless responses to build theory on and obtain better understanding of the nature of (partially) careless responding (cf. Gibbert et al.,, 2021, for theory-building based on outlying data points). Finally, we provide the freely available software package carelessonset (Welz & Alfons,, 2024), which implements CODERS in the statistical programming environment R (R Core Team,, 2024) can be downloaded from https://github.com/mwelz/carelessonset. As such, it is a useful and accessible tool for any researcher who is concerned with survey fatigue or careless responding in general.
Introducing CODERS
The literature generally identifies three distinct ways in which careless responding may manifest: low internal consistency of the given responses (inconsistency), low variability of given responses (invariability), and impossibly fast responses (e.g., Curran,, 2016; DeSimone et al.,, 2015; DeSimone & Harms,, 2018; Edwards,, 2019; Ward & Meade,, 2023). We expect partial careless responding to result in a change in respondent behavior in at least one of these dimensions. CODERS therefore offers a general framework for searching for a joint changepoint in any combination of measurements of these dimensions of carelessness. While the majority of this paper is focused on inconsistency and invariability, we revisit response times in a later section.
We propose to measure the first dimension, inconsistency, by means of reconstructions of observed responses (cf. Alfons & Welz,, 2024). These reconstructions are generated by an auto-associative neural network (henceforth autoencoder) (Kramer,, 1992), which learns response patterns that characterize attentive responding. We expect that random-like, content-independent responses cannot be learned well and are thus poorly reconstructed by the autoencoder, leading to a changepoint in reconstruction performance.
For measuring the second dimension, invariability, we propose a novel algorithm that is inspired by the famous longstring index of Johnson, (2005). Since long sequences of identical responses or constant response patterns are not expected in surveys that use positively and negatively coded items in a randomized order, we expect a changepoint in response variability once a respondent commences to respond carelessly through straightlining or pattern responding behavior.
Overall, CODERS attempts to capture evidence for the onset of carelessness by combining multiple indicators that are potentially indicative of such an onset, where the different indicators may capture different manifestations of carelessness. Conversely, if a respondent never becomes careless, we do not expect a changepoint in any dimension. Combining multiple indicators is a generally recommended practice to capture various types of careless responding (Goldammer et al.,, 2020; Huang et al.,, 2012; Meade & Craig,, 2012; Ward & Meade,, 2023).
We provide a detailed description of our assumptions in Appendix A and a technical definition of CODERS in Appendix B.
Quantifying Inconsistency with Autoencoders
Ward & Meade, (2023) define internal consistency of responses as patterns that are expected based on theoretical/logical grounds or trends in the data. For instance, items that are part of the same construct are expected to correlate strongly in most participants, provided that participants are attentive. In contrast, inconsistent careless responding “generate[s] responses that fail to meet an expected level of consistency” (Ward & Meade,, 2023). Respondents may engage in inconsistent careless responding if they attempt to conceal their carelessness, for instance by randomly choosing from all response options (Ward & Meade,, 2023). For our purposes, we consider the defining characteristic of inconsistent careless responding to be content-independent responses that are approximately randomly chosen from the response options, where the probability of being chosen may differ between response options, such as preferring extreme response options.
In order to identify inconsistent careless responding, we propose to use the machine learning method of autoencoders (Kramer,, 1992), which were originally developed to filter random noise in signal processing applications. An autoencoder is a neural network (e.g., Bishop,, 2006; Goodfellow et al.,, 2016) that attempts to reconstruct its input. In other words, the output variables are equal to the input variables. The idea behind reconstructing the input data is to learn the internal structures of the data, thereby separating the signal from random noise that lacks structure. Since we consider the defining characteristic of inconsistent careless responding to be near-randomly chosen responses—which may be viewed as random noise in the context of questionnaire data—we expect autoencoders to perform well in distinguishing between attentive responses and inconsistent careless responses.
To achieve the goal of learning the internal structures of a dataset, it is crucial to avoid that an autoencoder simply passes the input data through its network architecture. Consequently, autoencoders first express the input data in terms of a low-dimensional representation before reconstructing the input data again from this low-dimensional representation (cf. Kramer,, 1992, and Chapter 14.6 in Goodfellow et al.,, 2016). The use of such a low-dimensional representation makes autoencoders appealing for questionnaire data, as questionnaires typically contain multiple items to measure each construct of interest. In this context, the autoencoder aims to learn the patterns between the latent constructs in the low-dimensional space. For attentive responses, randomness that can be attributed to taking multiple measurements of the latent constructs is thereby filtered out. Since the amount of randomness from these multiple measurements should be relatively low for a well-designed questionnaire, attentive responses are expected to be reconstructed well from the learned patterns in the low-dimensional space of the latent constructs. Inconsistent careless responses, on the other hand, should contain little to no structure also in this low-dimensional space, and are therefore expected to be reconstructed poorly.
A schematic example of the network architecture of an autoencoder is given in Figure 1. Concretely, an autoencoder is a fully connected network whose nodes are organized in multiple layers. Each node takes a linear combination of the nodes from the previous layer and subsequently applies a so-called activation function in order to allow for nonlinear relationships. A crucial part of an autoencoder is the central layer, known as the bottleneck layer, which stores the low-dimensional representations of the input data. The network architecture of an autoencoder is symmetric around this bottleneck layer such that the can be viewed as having an encoder-decoder architecture: the encoder compresses the data into the low-dimensional representations, while the decoder reconstructs the input data again. It follows that an autoencoder can be seen as a dimension reduction technique, specifically as a nonlinear generalization of principal component analysis (PCA; see Kramer,, 1991).
Figure 1: Schematic example of a neural network with an autoencoder structure.
$\bm{x}_{1}$$\bm{x}_{2}$$\bm{x}_{3}$$\bm{x}_{4}$$\bm{\widehat{z}}_{1}$$\bm{\widehat{z}}_{2}$$\bm{\widehat{x}}_{1}$$\bm{\widehat{x}}_{2}$$\bm{\widehat{x}}_{3}$$\bm{\widehat{x}}_{4}$nput LayerMapping LayerBottleneck LayerDemapping LayerOutput Layer
Note. The network consists of five fully connected layers and is symmetric around the central layer. Each dimension is represented by one node, and nodes within the same layer are aligned vertically. In each node, a so-called activation function is applied to a linear combination of the nodes from the previous layer. The first layer (input layer) represents the input variables (here $\bm{x}_{1},\dots,\bm{x}_{4}$). The second layer (mapping layer) prepares the data to be compressed in the third and central layer (bottleneck layer), which holds low-dimensional representations of the input data (here $\bm{\widehat{z}}_{1}$ and $\bm{\widehat{z}}_{2}$). The fourth layer (demapping layer) prepares the low-dimensional respresentations for reconstructing the input data in the final layer (output layer), yielding reconstructions in the original dimension (here $\bm{\widehat{x}}_{1},\dots,\bm{\widehat{x}}_{4}$).
Hence, the number of nodes in the bottleneck layer is an important design choice in the architecture of an autoencoder. In the context of questionnaire data, we recommend setting this number equal to the number of constructs measured in the questionnaire. For example, consider data from the Revised NEO Personality Inventory (NEO-PI-R; Costa & McCrae,, 1992), which contains 240 items and measures six so-called facets of each of the Big Five personality traits. Thus, this instrument measures a total of 30 facets, such as anxiety and modesty, so that we would set the number of nodes in the bottleneck layer to 30.
A thorough mathematical definition of autoencoders can be found in Appendix B.1. Although there are other design choices besides the number of nodes in the bottleneck layer (see Appendix B.1.2), our software provides default values that we recommend researchers to use. The only input required by the user is the number of constructs measured in the survey (for the number of nodes in the bottleneck layer).
Suppose now that we have obtained autoencoder reconstructions of the responses of each survey participant. Let there be $n$ participants and $p$ items, and denote the observed response of the $i$-th participant to the $j$-th item by $x_{ij}$ and the corresponding reconstruction by $\widehat{x}_{ij}$, for $i=1,\dots,n$ and $j=1,\dots,p$. We define the reconstruction error $\textsf{RE}_{ij}$ associated with the response of participant $i$ to item $j$ to be the squared difference between the observed response and its reconstruction, scaled by the range of the answer categories. Formally, for participants $i=1,\dots,n$ and items $j=1,\dots,p$,
$\textsf{RE}_{ij}=\left(\frac{x_{ij}-\widehat{x}_{ij}}{L_{j}-1}\right)^{2},$ (1)
where $L_{j}$ denotes the number of answer categories of item $j$. Recall that we expect the reconstruction errors $\textsf{RE}_{ij}$ to be low in case of attentive responding but high in the presence of inconsistent careless responding. Hence, if participant $i$ starts providing inconsistent careless responses at item $k\in\{1,\dots,p\}$, we expect a changepoint at position $k$ in the participant’s series of $p$ reconstruction errors, as reconstruction errors are expected to be higher from item $k$ onward. An example can be found in the top panel of Figure 2(a).
Figure 2: Illustrative example output from CODERS using responses of a participant to a 300-item personality inventory.
(a) Two dimensions indicative of carelessness: autoencoder reconstruction errors (RE) in the top panel, and the LongstringPattern (LSP) sequence in the bottom panel.
(b) The per-response statistic of the changepoint detection test. The dashed horizontal line corresponds to the associated critical value at an extremely conservative significance level of 0.1%.
Note. The solid vertical lines indicate the identified carelessness onset, which is at the 148th item. Reconstruction errors are in general much larger after this item than before, indicating more random-like responses. The response data are taken from Johnson, (2005) and are discussed in detail in a later section.
Quantifying Invariability
We propose to measure response invariability through a novel algorithm inspired by the longstring index of Johnson, (2005) that we call LongstringPattern (LSP). This algorithm exploits that invariable careless responses are characterized by content-independent response patterns. A pattern sequence has the defining property that it consists of repeated occurrences of the same response pattern. Consider the sequence of responses $-2-1-2-\ldots-1-$ 2 - … - 1 - 2, which consists of repeated occurrences of the pattern $1-2$. Denoting the length of a pattern by $l$, the above example exhibits a pattern of length $l=2$. Analogously, the sequence $-3-1-2-3-\ldots-1-2-$ 3 - … - 1 - 2 - 3 consists of repeated occurrences of the pattern $1-2-3$ of length $l=3$, whereas the straightlining sequence $1-1-\ldots-$ 1 - … - 1 consists of repeated occurrences of 1, which is a pattern of length $l=1$.
Define $l\textsf{-pattern}_{ij}$ as the number of consecutive items in the longest sequence of a recurring pattern of length $l$ in which the response of participant $i$ to item $j$ takes part, or assign the value 1 if the response is not part of a recurring pattern of length $l$. Consider the following two illustrative examples. First, for $l=2$, the response sequence $1-2-1-2-1-4-3-5-4-5-4$ will be assigned the 2-pattern sequence $4-4-4-4-1-1-1-4-4-4-4$ because the first four responses and last four responses comprise two occurrences of the distinct patterns $1-2$ and $5-4$, respectively, which are both of length $l=2$. The central three responses ($1-4-3$) are not part of any recurring pattern of length $l=2$, hence they are each assigned the value 1. Second, for a recurring pattern of length $l=1$ (i.e., consecutive identical responses), the response sequence $3-2-3-3-1-4-1-1-1$ is assigned the 1-pattern sequence $1-1-2-2-1-1-3-3-3$, since the subsequences of consecutive identical responses $3-3$ and $1-1-$ 1 - 1 comprise two and three responses, respectively.33endnote: 3The 1-pattern sequence can be seen as a generalization of the the longstring index of Johnson, (2005), which we recover by calculating $\max_{1\leq j\leq p}1\textsf{-pattern}_{ij}$ for each respondent $i=1,\dots,n$. In short, higher values in the $l$-pattern sequence capture a tendency to respond in fixed patterns of length $l$. If a participant starts engaging in invariable careless responding, we expect a changepoint in the $l$-pattern sequence (for some $l$) from relatively low to relatively high values, assuming that the order of items is randomized.
However, an $l$-pattern sequence crucially depends on the choice of pattern length $l$, which is restrictive since invariable careless respondents may choose widely different patterns. Consequently, a single choice of pattern length $l$ is unlikely to capture all careless response patterns that may emerge. To tackle this issue, we propose to calculate the $l$-pattern sequence for multiple choices of $l$, and we define
$\textsf{LSP}_{ij}=\max_{l}\ l\textsf{-pattern}_{ij}$ (2)
for participant $i=1,\dots,n$ and item $j=1,\dots,p$. In this paper, we vary $l$ from 1 to the number of answer categories in the survey. The rationale behind this choice is that we consider it unlikely that careless respondents choose complicated individual patterns whose length exceeds the number of answer categories.
We provide a detailed description of the LSP algorithm in Appendix B.2. An example of an LSP sequence is shown in the bottom panel of Figure 2(a).
Identifying Carelessness Onset via Changepoint Detection
For each participant, we can obtain multiple individual series of length equal to the total number of items in the survey. Each individual series should thereby measure one of the dimensions indicative of carelessness. Although we focus on using reconstruction errors from (1) to measure inconsistency as well as the LSP sequence from (2) to measure invariability, it should be noted that in principle CODERS can also be used with any single measurement or any other combination of different measurements of carelessness.
As motivated in the previous two sections, we expect a changepoint in any of the measurements of carelessness from the onset of careless responding (e.g., Figure 2(a)), and no changepoint in the absence of carelessness. We therefore propose to apply the nonparametric cumulative sum self-normalization test of Shao & Zhang, (2010), which was developed in the context of time series analysis. It searches for the location of a possible changepoint in a multivariate series and has attractive theoretical guarantees (see Shao & Zhang,, 2010; Zhao et al.,, 2022, and our detailed description in Appendix B.3).
For each point in a multivariate series (we use measurements of carelessness associated with each questionnaire item), this changepoint detection test calculates a statistic that compares a quantity of interest (we take the mean) in the periods before and after the current point (item). If the maximum of this statistic exceeds a critical value, the point (item) where this maximum occurs is flagged as a changepoint. Otherwise no changepoint is flagged. The critical value thereby depends only on the significance level $\alpha$ and the dimension of the multivariate series. Importantly, it controls the size of the test regardless of the distribution of the individual series. This gives CODERS great flexibility, as it allows us to use any combination of measurements of carelessness without making possibly unrealistic assumptions on their distributions or obtaining cutoff values in an ad-hoc manner. In line with literature that suggests that there should be overwhelming evidence when labeling respondents as careless (e.g., Huang et al.,, 2012), we recommend using a conservative significance level no greater than 1% for identifying changepoints.
As an example, consider again the two-dimensional series in Figure 2(a), which measures carelessness of a certain survey participant in terms of autoencoder reconstruction errors (RE) and the LongstringPattern (LSP) sequence. Figure 2(b) shows the corresponding series of statistics, where the maximum value at the 148th item clearly exceeds the critical value. That is, CODERS flags this respondent as being careless from the 148th item onwards. Indeed, from roughly this item onwards, the reconstruction errors are in general much larger than before. It follows that this respondent likely is a partially inconsistent careless respondent.
Empirical Application
Ideally, new methodology is empirically validated by conducting an experiment that emulates the phenomenon of interest in lab conditions. For instance, Schroeders et al., (2022) conduct an experiment in which they instruct one part of the participants to respond truthfully and another part to respond carelessly in order to investigate whether various screeners accurately identify the subset of careless respondents. For any of the compared methods, at most 19% of the respondents flagged as careless were in fact instructed to be careless. The authors therefore express doubts whether the study participants complied with the experimental instructions. In addition to the risk of non-compliance, Ulitzsch et al., 2022a ; Ulitzsch et al., 2022b stress that another major limitation of this experimental approach is the doubtful assumption that participants instructed to respond carelessly behave in a comparable manner to careless respondents outside of such lab experiments. Indeed, it seems plausible that careless responding stems from an intrinsic state-of-mind rather than being a conscious action that can be instructed (cf. Alfons & Welz,, 2024, for additional discussion). Hence, rather than conducting such an experiment, we take data from a seminal study on careless responding and further examine the responses for partial carelessness.
Data
We apply CODERS to an empirical data set collected by Johnson, (2005). The data concern the five factor model of personality (Goldberg,, 1992), which assumes that variation in a measurement of personality can be explained by the so-called Big Five personality traits (factors): openness, conscientiousness, extraversion, agreeableness, and neuroticism. Each of the Big Five factors can be further decomposed into six subdomains (facets): for instance, depression and anxiety are facets of the neuroticism trait. Consquently, common Big Five instruments, such as the 240-item NEO-PI-R (Costa & McCrae,, 1992), typically measure a total of $5\times 6=30$ facets. The data set of Johnson, (2005) consists of internet-collected responses to a 300-item representation of the NEO-PI-R inventory (Goldberg,, 1999), in which each of the 30 facets is measured by 10 items on a five-point Likert scale.
[…]
In general, it seems likely that survey fatigue is present in a sample of responses to an online questionnaire with a lengthy battery of 300 items. For instance, Bowling et al., (2021) estimate that if one wishes that at least 90% of participants do not respond carelessly to more than 20% of all items, one should not include more than 176 items in an online questionnaire. For the study of Johnson, (2005), it is important to keep in mind that the author applied two methods for detecting inconsistent careless respondents to the data from which straightliners are already removed. The two methods are designed to detect participants who respond carelessly to all or at least the vast majority of survey items. It is unclear how well they perform in identifying partially careless respondents, who may respond carelessly to only a relatively small number of items. In addition, Johnson, (2005) did not apply any method designed to identify invariable careless responding other than straightlining. By applying CODERS, which is designed to identify partially careless responding for various types of carelessness, new insights into the incidence and nature of (partially) careless responding may emerge.
Results
[…]
Comparison with Benchmark Methods
For obtaining the autoencoder reconstruction errors in CODERS, we set the number of nodes in the bottleneck layer to 30, i.e., equal the number of facets in the Big Five model. Table 1 summarizes the results of CODERS. Depending on the significance level $\alpha$, CODERS flags between 164 (0.7%) participants for $\alpha=0.001$ and 850 (3.8%) participants for $\alpha=0.01$ as partially careless. On average, the location of the carelessness onset item for flagged participants is estimated around the 150th item, meaning that carelessness onsets on average after half of the 300 total items. These findings are broadly in line with the results of Bowling et al., (2021, Figure 1), whose model suggests a low probability that a representative participant starts responding carelessly at the 150th item. The earliest onset occurs at item 12, whereas the latest onset is at item 294. A more detailed visual overview of the detected onset items for the flagged participants can be found in Figure C.1 in Appendix C.2.
For comparison, we consider personal reliability and psychometric antonym scores. Along the given responses, Johnson, (2005) provides per-individual estimates for either score, which we use in our analysis. Importantly, Johnson, reverses psychometric antonym scores so that low scores are associated with carelessness in both methods. Following Johnson,, we flag respondents as careless whose personal reliability and psychometric antonym scores are lower than $0.3$ and $-0.03$, respectively. The results are shown in Table 1. Personal reliability flags 40 participants (0.2%) as careless, while psychometric antonym flags 186 respondents (0.8%). CODERS, personal reliability, and psychometric antonyms generally show little agreement regarding the detected careless respondents. Personal reliability and psychometric antonym scores correlate by a moderate value of 0.488, yielding only 13 participant that are flagged by both methods. Depending on the significance level $\alpha$, CODERS and personal reliability have between 1 and 4 flagged participants in common, while the number of jointly flagged respondents for CODERS and psychometric antonym ranges between 5 and 10. All three methods share a single flagged participant.
Table 1: Comparison of CODERS with benchmark methods for the data of Johnson, (2005).
Also flagged by CODERS
Method Number of flagged respondents $\alpha=0.001$ $\alpha=0.005$ $\alpha=0.01$
Reliability 40 (0.2%) 1 3 4
Antonym 186 (0.8%) 5 7 10
Both 13 1 1 1
[…]
It is not surprising that there is little to no consensus of CODERS with the two benchmark methods, as the former is designed for a different task: finding respondents that have been careless in parts of the questionnaire (including identifying the onset item) rather than throughout the entire questionnaire. This disagreement indicates that partial carelessness may be present in the data, which the benchmark methods may be unable to identify. We therefore take a closer look at a representative selection of respondents and analyze their per-item reconstruction errors and LSP sequences, which are shown in Figure 3. Table 2 contains additional summary statistics for these measurements before and after the estimated careless onset item (if such an item is found by CODERS). Two of the selected participants are flagged as careless by CODERS but not by either of the two benchmark methods, one participant is flagged by both CODERS and psychometric antonym, and one participant is flagged by both benchmark methods but not CODERS.
Figure 3: Scores and careless onset item detected by CODERS in selected participants from the data of Johnson, (2005).
(a) Flagged by CODERS only, presumably inconsistent. Onset is estimated at item 148.
(b) Flagged by CODERS only, presumably invariable. Onset is estimated at item 58.
(c) Flagged by antonym and CODERS, presumably inconsistent. CODERS estimates onset at item 119.
(d) Flagged by antonym and personal reliability, but not CODERS (i.e., there is no estimated onset item).
Note. In each subplot, autoencoder reconstruction errors (RE) and LongstringPattern (LSP) sequences are shown in separate panels. The solid vertical lines indicate the identified careless onset items. Summary statistics for the corresponding participants are provided in Table 2.
Table 2: Average scores before and after the careless onset item detected by CODERS in selected participants from the data of Johnson, (2005).
Before After
RE 0.019 (0.035) 0.069 (0.082)
LSP 4.197 (1.396) 3.497 (1.433)
(a) Flagged by CODERS only, presumably inconsistent. Onset is estimated at item 148.
Before After
RE 0.078 (0.103) 0.025 (0.035)
LSP 2.421 (0.706) 19.770 (10.941)
(b) Flagged by CODERS only, presumably invariable. Onset is estimated at item 58.
Before After
RE 0.019 (0.022) 0.039 (0.059)
LSP 3.619 (1.371) 2.687 (0.972)
(c) Flagged by antonym and CODERS, presumably inconsistent. CODERS estimates onset at item 119.
Overall
RE 0.064 (0.066)
LSP 2.660 (0.753)
(d) Flagged by antonym and personal reliability, but not CODERS (i.e., there is no estimated onset item).
Note. Standard deviations are given in parentheses. The subtables correspond to the four selected participants from Figure 3. Positioning of subtables is consistent with positioning of the subfigures in Figure 3, e.g., Table 2(a) refers to the same participant as Figure 3(a).
First, we focus on Figure 3(a) and Table 2(a). CODERS identifies carelessness onset at item 148, but neither benchmark method flags this respondent as careless. We observe that the average reconstruction error, as well as the variability in the reconstruction errors, is much larger before the estimated onset item than afterward. Since the pre-onset responses could be reconstructed well (except for a brief spike around item 110) while post-onset responses could not, there is evidence that the participant started the questionnaire attentively, but succumbed to survey fatigue about halfway through and subsequently engaged in inconsistent careless respondening.
The second example from Figure 3(b) and Table 2(b) shows another participant that is flagged by CODERS but no benchmark method. The estimated onset of careless responding is at item 58. While the LSP sequence clearly changes around this item, even more excessive values are found for later items, with a maximum LSP value of 31. In fact, only four out of the $n=20,994$ participants in the sample yield more extreme LSP values. Hence, there is strong evidence that this participant responded with (increasingly) lengthy sequences of fixed patterns as the questionnaire progressed. That is, they likely started engaging in invariable careless responding.
The third example, shown in Figure 3(c) and Table 2(c), is similar to the first example in the sense that it suggests an initially attentive respondent starting to give inconsistent careless responses from some point onward in the questionnaire, as indicated by a higher average and a larger variability in the reconstruction errors. Here, the careless onset is estimated at item 119. The crucial difference with the first example is that this participant was not only flagged as careless by CODERS, but also by psychometric antonym. A plausible explanation for this difference may be found in the earlier careless onset. If a participant responds carelessly to a large enough majority of the items, a method designed to detect carelessness throughout the questionnaire (such as the benchmark methods) has a better chance at flagging that participant as careless.
[…]
While these four examples are naturally selective, similar plots as in Figure 3 for all participants flagged by CODERS or the two benchmark methods can be found in our online repository at https://github.com/mwelz/carelessonset_plots. In general, two key insights can be derived from the above analysis. First, we find empirical evidence that CODERS can indeed detect various types of partial careless responding, specifically inconsistent and invariable carelessness. Common screeners for careless responding throughout the questionnaire, such as personal reliability or psychometric antonym, are typically only effective for one type of careless responding. Second, CODERS is not a replacement for existing screeners, but rather a complimentary tool that provides new insights if participants display careless responding behavior only in part of the questionnaire.
Comparison with Johnson, (2005)
We now compare our findings with those reported by Johnson, (2005) regarding the incidence of careless responding. This comparison is summarized in Table 3.
Table 3: Incidence of careless responding in the data of Johnson, (2005).
Johnson, (2005) Our findings
Type of careless responding Method Number Method Number
Straightliners Longstring + screeplot 808 (3.6%)
Longstring + cutoff by Costa & McCrae, (2008) 1455 (6.5%)
Inconsistent Reliability 41 (0.2%) Reliability 40 (0.2%)
Antonym 185 (0.8%) Antonym 186 (0.8%)
Either 213 (1.0%) Either 213 (1.0%)
Partial CODERS ($\alpha=0.001$) 158 (0.7%)
CODERS ($\alpha=0.005$) 580 (2.6%)
CODERS ($\alpha=0.010$) 836 (3.7%)
[…]
Johnson, explores two different cutoff values for the longstring index to identify straightliners: one obtained via a screeplot and another suggested by Costa & McCrae, (2008),66endnote: 6Johnson, (2005) refers to a pre-publication version of Costa & McCrae, (2008). with the latter yielding almost twice as many flagged respondents (3.6% vs. 6.5%). As the data set provided by the author already excludes straightliners (among others), we could not attempt to replicate these findings.
Regarding inconsistent careless responding, we find minor differences in the number of flagged respondents compared to Johnson,: 40 vs. 41 for personal reliability, and 186 vs. 185 for psychometric antonym. These minor differences are potentially due to rounding: the psychometric antonym and personal reliability scores provided by Johnson, are rounded by two decimal places, whereas the scores used to compute the reported results in Johnson, (2005) were likely unrounded. As in Johnson, (2005), we find that personal reliability and psychometric antonym agree on 13 of the flagged participants.
Following strategies that “[eliminate] more non-attentive responders, albeit at the cost of more false positives”, Johnson, (2005, p. 116) applies the cutoff value of Costa & McCrae, (2008) for identifying straightliners, and considers respondents that are found by either personal reliability or psychometric antonym as inconsistent careless respondents. This results in a total of 1668 detected careless respondents (7.4%). Employing a similar strategy regarding CODERS, one may choose a significance level of 1%. Adding the corresponding 836 participants yields a total of 2504 flagged careless respondents (11.2%).
Hence, we find evidence that the incidence of careless responding might be substantially underreported in Johnson, (2005) due to hitherto undetected partially careless respondents. Moreover, an incidence of 11.2% in this data set would be higher than the 5–10% that the literature deems problematic for further analysis (e.g., Arias et al.,, 2020; Credé,, 2010).
Simulation Experiments
After demonstrating how CODERS yields new insights in data from a seminal study, we further study the behavior of CODERS via simulation experiments inspired by existing survey measurements.
Data Generation
We generate rating-scale responses of $n=500$ respondents to $p=300$ items. Each item has five Likert-type answer categories (anchored by 1 = “strongly disagree” and 5 = “strongly agree”). The simulated questionnaire measures 30 constructs, each of which is measured by a scale of 10 items. We simulate reliable scales by imposing that items from the same scale have a correlation of 0.7, which results in a Cronbach’s $\alpha$ value of 0.959.77endnote: 7We use package simstudy (Goldfeld & Wujciak-Jens,, 2020) to generate the data, for which the target values of correlations only hold approximately on the population level. Hence, the same applies to Cronbach’s $\alpha$ values. Items from different scales are uncorrelated. All 10 items within a given scale follow the same response probability distribution. We thereby consider the two types of distributions defined in Table 4: centered about the midpoint and skewed towards agreeing. Each of the two distributions is used in 15 scales, yielding the total of $2\times 15\times 10=300$ items. We reverse five items per scale (i.e., half the items) and randomize the order of the items. Note that distributions skewed towards disagreement are implicitly covered by this design because of the item reversal.
Table 4: Response probability distributions of items within different scales.
Type of distribution $P[X=1]$ $P[X=2]$ $P[X=3]$ $P[X=4]$ $P[X=5]$
Centered 0.15 0.20 0.30 0.20 0.15
Skewed 0.10 0.15 0.20 0.25 0.30
Note. The random variable $X$ denotes the response to an item with five Likert-type answer categories.
Subsequently, we select a fixed percentage $\gamma\in\{0,0.02,0.04,0.1,0.2,\dots,0.6,0.8,1\}$ of respondents to be partially careless. That is, $\gamma$ denotes the prevalence of careless responding. For each of the selected partially careless respondents, we sample a carelessness onset item from which onward the remaining responses are careless. We consider two settings for the carelessness onset: (i) early onset, which we randomly sample from the item set $\{30,31,\dots,269,270\}$ so that careless responding starts between 10% and 90% of all items, and (ii) later onset, which we randomly sample from $\{150,151,\dots,269,270\}$, i.e., between half and 90% of all items.
[…]
We repeat the above described data generating process 100 times. Importantly, we keep the location of carelessness onset item of each participant constant across the 100 repetitions, which allows for neat visualization of the results. We compare three variations of CODERS based on (i) both reconstruction errors and the LSP sequence, (ii) reconstruction errors only, and (iii) the LSP sequence only. For the autoencoder, we set the number of nodes in the bottleneck layer equal to the number of measured constructs (i.e., 30).
Performance Evaluation
To quantify how accurately the onset of careless responding is detected, we compute the adjusted Rand index (ARI) (Hubert & Arabie,, 1985, see Appendix D.1 for details). The ARI is best known as a measure of clustering performance: a value of 0 is expected for a random clustering, while a perfectly recovered clustering results in a value of 1. In the careless onset detection problem for a given respondent, we can view the items before the careless onset as one cluster and the items from the onset onwards as another. Specifically, the closer the ARI is to its maximum value 1, the better CODERS performs at accurately locating the true carelessness onset item. On the other hand, an ARI value close to 0 indicates that the estimated onset is far from the true onset. If a respondent is attentive throughout the questionnaire but CODERS incorrectly identifies a changepoint, the ARI takes a value of 0. The same holds if CODERS fails to flag a changepoint for a partially careless respondent.
We report the average ARI across respondents. Note that averaging the ARI over attentive respondents yields the proportion of attentive respondents who are correctly identified as attentive. Conversely, subtracting this number from 1 gives the proportion of attentive respondents that are incorrectly flagged as partially careless.
Results
We start presenting the results of CODERS for the three dimensions by means of a visual inspection. For a carelessness prevalence level of $\gamma=0.6$ and the early onset setting, Figure 4 displays a heatmap holding the indices of the participants in its rows and the item indices in its columns. For participant $i\in\{1,\dots,n\}$ and item $j\in\{1,\dots,p\}$, the color scale of the cell at position $(i,j)$ measures the frequency across the 100 repetitions with which the response of participant $i$ to item $j$ has been flagged as changepoint: a more saturated red indicates that the corresponding item was flagged more frequently as a changepoint. The black rectangles visualize the true location of the careless onset items (recall that those are fixed across the repetitions). The $y$-axis of the plot is rearranged so that the participant indices are ordered according to their respective carelessness onset. Good detection performance is characterized by flagging cells that are close to the corresponding onset items. Indeed, we can see that there is a saturated red band that closely follows the black rectangles. Note that the top 40% of the rows in Figure 4 very rarely see flagged changepoints. Those are the respondents who never become careless. Based on this visual examination, CODERS seems to not only accurately estimate the careless onset, but also discriminate well between partially careless respondents and fully attentive respondents. Results are similar for other combinations of carelessness prevalence and onset, hence we omit the corresponding heatmaps.
Figure 4: Heatmap of carelessness onset detection performance of CODERS.
Note. Of the $n=500$ respondents, 60% are partially careless according to the early onset setting. The significance level of CODERS is set to 0.1%. A more saturated red in a cell corresponds to a higher frequency across the 100 repetitions of the cell being flagged as carelessness onset item ($x$-axis) of a given respondent ($y$-axis). The black rectangles denote the true carelessness onset items (which are constant across the repetitions) and the $y$-axis is rearranged according to the location of true carelessness onset. An absence of a black rectangle (occurs in 40% of all respondents) indicates that a given respondent never becomes careless.
Figure 5: ARI of CODERS for various carelessness prevalence levels.
Note. The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows), and the significance level of CODERS is set to 0.1%.
Varying the prevalence of careless responding, Figure 5 visualizes the ARI averaged over attentive respondents (left column) and partially careless respondents (right column) for the early onset (top row) and later onset (bottom row) settings. Each of the three variations of CODERS rarely flag changepoints in fully attentive respondents, except when carelessness prevalence is extremely high. While the percentage of false positives is close to the significance level 0.1% until 40% prevalence, overrejection remains moderate at 1% for 60% prevalence, before it rises to about 5% for early onset and 10% for later onset at 80% prevalence. For partially careless respondents, the ARI of CODERS with both dimensions remains remarkably constant throughout carelessness prevalence levels at about 0.75 and 0.85 for early and later onset, respectively, mirroring an excellent performance of flagging changepoints in correct locations. In contrast, the ARI using only the LSP sequences amounts to about 0.6 throughout prevalence levels in both onset settings, whereas the ARI using only the reconstruction errors drops from about 0.5 for 2% prevalence to about 0.3 for 100% prevalence with early onset, and from about 0.75 to about 0.35 with later onset. It is important to note that such a drop in performance is expected when using only one indicator of carelessness due to various types of careless responding being simulated. Reconstruction errors are designed to capture inconsistent carelessness and may therefore miss invariant carelessness, while LSP sequences are designed to capture invariant carelessness and may therefore miss inconsistent carelessness.
Figure 6: ARI of CODERS for various carelessness prevalence levels.
Note. The ARI is averaged over the participants corresponding to each careless response type (columns) for different carelessness onset settings (rows), and the significance level of CODERS is set to 0.1%. Note that for straightlining and pattern responding, the results for CODERS using all two dimensions and LongstringPattern only are nearly identical, resulting in overplotted lines.
To explore this further, Figure 6 displays the ARI averaged over subgroups of participants corresponding to each type of careless responding (separate columns) for early onset (top row) and later onset (bottom row). Indeed, reconstruction errors yield much better performance for inconsistent carelessness (random and extreme responding) than for invariable carelessness (straightlining and pattern responding), in particular in the later onset setting. Conversely, using LSP sequences results in near-perfect detection of straightlining and pattern responding with an ARI of about 0.99, but performs particularly poorly for random responding with an ARI close to 0. This highlights the importance of combining the two dimensions in CODERS for accurately detecting various types of partially careless responding.
We now zoom in on the results in Figure 6 for CODERS using reconstruction errors only, which vary by type of (inconsistent) carelessness, as well as by carelessness prevalence and onset location. First, the performance is clearly better for extreme responding than for random responding, with the ARI being in general higher and remaining more stable across prevalence levels. Considering the relatively flat probability distributions of the attentive responses (see Table 4), distinguishing random responses is simply a much more challenging problem. Second, the heterogeneity in performance across prevalence levels and onset location is likely due to the autoencoder requiring sufficiently many attentive responses to successfully distinguish between attentive and (inconsistent) careless responding. If carelessness tends to onset later in the survey, the autoencoder has more information to learn attentive response patterns, making it easier to distinguish between attentive and careless responses. Similar reasoning applies for carelessness prevalence. When fewer responses are careless, the autoencoder has access to more attentive responses from which to learn. In general, these results suggest that CODERS can accurately identify inconsistent careless responding unless careless responses are excessively present in the data through a high prevalence of careless respondents whose carelessness onset occurs early in the survey. However, having an overwhelmingly large number of participants with largely uninformative responses suggests a fundamental issue with data quality, in case of which any subsequent analysis may yield questionable results.
To summarize, this simulation experiment suggests that CODERS succeeds in combining complimentary indicators for carelessness onset in order to capture different manifestations of careless responding.
Additional Simulations
We have conducted a variety of additional simulation experiments, which are described and discussed in Appendix D.2. Specifically, we investigate different choices of the significance level in CODERS, varying correlation structures (and therefore measurement reliability), as well as varying sample sizes. Overall, the findings presented above are representative for those situations. For low to moderate carelessness prevalence (as in the empirical application), choosing a higher significance level (e.g., 1%) leads to an increase in power while retaining size control. On the other hand, for high prevalence levels (about 40% or higher), increasing the significance level yields too many false positives. As expected, power decreases (increases) slightly with decreasing (increasing) measurement reliability or sample size.
We conclude from our simulation experiments that CODERS reliably detects carelessness onset while successfully discriminating between partially careless respondents and fully attentive respondents. In addition, our simulations highlight the importance of using multiple indicators to capture different manifestations of carelessness (cf. Goldammer et al.,, 2020; Huang et al.,, 2012; Meade & Craig,, 2012; Ward & Meade,, 2023).
Extensions, Limitations and Alternatives
Below, we address the potential incorporation of response times into CODERS, and we discuss general limitations as well as potential alternatives.
Response Times Revisited
In addition to inconsistency and invariability, fast responding is considered the third major manifestation of carelessness (e.g., Ward & Meade,, 2023). In (online) surveys, response time is typically measured by the total time a participant spent on the questionnaire, time spent on each questionnaire page, or (less common) time spent on each questionnaire item.
If per-item response times are available, they can easily be incorporated into CODERS as an additional series to be used in the changepoint detection. If only per-page response times are available and pages are relatively short, an additional series can be obtained by assigning to each response the average page time, i.e., the time the participant spent on the page on which the item is located divided by the number of items on that page (which inspired by the page time index of Bowling et al.,, 2023). If a participant at some point in the survey, for instance, starts responding without reading the items, we expect a changepoint in response time towards considerably faster responses (cf. Bowling et al.,, 2023, 2016; Meade & Craig,, 2012; Huang et al.,, 2012).
Nevertheless, it is unclear if a changepoint in response times is indeed indicative of careless responding in an empirical setting. This may depend on the survey design: if items in one part of the survey are sufficiently longer than in other parts, a changepoint in response times might occur naturally for attentive respondents. On the other hand, if all items are of similar length, a changepoint to (much) shorter response times may indicate carelessness. For instance, there exist instruments whose individual items comprise of single words, such as the Big Five unipolar markers (Goldberg,, 1992) with items like “timid” or “envious”. In such measurements, one would not expect large differences in response times across items. For such survey designs with similar item length, incorporating response times into CODERS may prove fruitful, but a thorough validation is beyond the scope of this paper and therefore left for future research.
Limitations
Throughout this paper, we have assumed that all participants start as truthful and accurate respondents, while some of them may begin to respond carelessly from a certain item onward. However, in certain scenarios, this assumption may not hold true.
First, as we have already seen in our empirical application, some participants may respond carelessly throughout all survey items (cf. Figure 3(d) and Table 2(d)). Consequently, CODERS is unlikely to flag a changepoint due to an absence of a change in behavior. Nevertheless, the literature has established a plethora of methods for detecting respondents who have been careless throughout the questionnaire. In this sense, CODERS—which is intended for detecting partial carelessness—is complementary to those existing methods. If one suspects that respondents who have been careless throughout all items are present, we recommend to apply established detection methods following the guidelines of Arthur et al., (2021) and Ward & Meade, (2023). Respondents that are flagged as careless by such methods, but for whom no changepoint was flagged by CODERS, have likely been careless throughout the entire survey.
Second, there may be some participants who alternate between periods of attentive and careless responding (Berry et al.,, 1992; Clark et al.,, 2003; Meade & Craig,, 2012). In this case, CODERS may only flag one of the possibly multiple changepoints. CODERS should therefore be viewed as a promising first step towards detecting periods of careless responding, with an extension for the detection of multiple changepoints being an area of further research.
Finally, the changepoint detection procedure of Shao & Zhang, (2010) employed by CODERS is based on asymptotic arguments when the number of items is large. Hence, CODERS is designed for lengthy questionnaires, and it is likely that the statistical power drops considerably in short questionnaires. However, partial careless responding due to fatigue may be less prevalent in short questionnaires (cf. Bowling et al.,, 2021). Studying the onset of careless responding in shorter surveys is therefore beyond the scope of this paper.
Potential Alternatives to CODERS
First, the recommendation that one should not include more than three detection items (Meade & Craig,, 2012) is fairly restrictive as carelessness may have onset at a much earlier item than the detection item, resulting in a failure to exclude a potentially large number of careless responses. In principle, a form of carelessness onset detection may be possible by a priori including detection items in the survey, such as bogus, instructed, or self-report items: one may consider a failed detection item (i.e., the participant provides a response that would not be expected from an attentive respondent) as an estimate of the onset of careless responding. However, there are a number of practical and theoretical disadvantages with this approach (see Curran,, 2016, pp. 13–15, for a more general discussion on the caveats of detection items). First, the estimates of the carelessness onset may be crude at best. Recommendations of Meade & Craig, (2012) suggest including at most three bogus, instructed, or self-report items, spaced every 50–100 items. Consequently, the carelessness onset may occur much earlier than at at the detection item, leaving a potentially large number of careless responses undetected. Second, attentive respondents may ironically choose an unexpected response to a bogus item because they find doing so humorous (Meade & Craig,, 2012) or due to misinterpretation of the item (Curran & Hauser,, 2019).88endnote: 8For instance, Curran & Hauser, (2019) find that some respondents would agree to the popular bogus item “All my friends say I would make a great poodle” because they believe that they are loyal friends and they associate loyalty with dogs. Third, a careless respondent may choose the expected response by mere chance or lie in self-report items (DeSimone & Harms,, 2018; Edwards,, 2019), resulting in failure to identify this careless respondent. Fourth, it is unclear if, e.g., the response “somewhat disagree” to a detection item for which “strongly disagree” is expected should be considered careless (Ward & Meade,, 2023). Finally, it may not be possible to include detection items in the survey for legal or ethical reasons, or because the survey analyst has no control over the survey design.
In addition, the software packages careless (Yentes & Wilhelm,, 2021) and ufs (Peters & Gruijters,, 2021) support the calculation of an intra-individual variance (IRV) across subsets of responses.99endnote: 9Specifically, this is implemented in the functions irv() from package careless and carelessReport() from package ufs. If a set of responses towards the end of the questionnaire exhibits a sufficiently larger (Marjanovic et al.,, 2015) or lower (Dunn et al.,, 2018) variance than an earlier set of responses, there might evidence for an onset of partially careless responding. First, these packages only support splitting the responses into subsets of equal length, which does not allow for accurate detection of the carelessness onset. Second, a more general implementation for confirming the presence of partial carelessness in arbitrary subsets would require a priori knowledge on the onset item that defines the subsets, which is unrealistic in practice. Third, it is unclear when a variance can be considered sufficiently low or high to infer carelessness: determining cutoff values below or above which a participant is to be flagged as careless may depend on survey design and survey content.1010endnote: 10For example, if 5-point scales are used a lower cutoff is necessary than for a 9-point scale, since IRVs will be very different for those scales.
Instead of screening for careless responding after data collection, one may attempt to prevent such behavior during survey administration. Preventive measures against careless responding may entail promising rewards for responding accurately (Gibson & Bowling,, 2020), as well as threatening with punishment for careless responding or the presence of a proctor (Bowling et al.,, 2021). There is evidence that each of these three preventive measures can prevent or at least delay carelessness onset (Bowling et al.,, 2021; Gibson & Bowling,, 2020; Huang et al.,, 2012). However, either of these three preventive measures can be undesirable or unethical in practice. Administering proctored questionnaires is often impractical and expensive, and is typically infeasible in online studies. Arthur et al., (2021) and Bowling et al., (2021) advise against using threats if one wishes to maintain a positive long-term relationship with questionnaire participants, such as when there is a professional relationship between researcher and participants. Moreover, threats seem inappropriate in longitudinal studies or when the participants are members of a questionnaire panel (as is common in survey research), or when they are suspected to suffer from anxiety, depression, or trauma (as is common, e.g., in research on wellbeing in the workplace, or more general in psychological research). On the other hand, promising rewards may cross ethical boundaries in organizational surveys and in the context of employment-related decisions (Arthur et al.,, 2021).
Notwithstanding, we stress that preventive methods against carelessness are complementary to detection methods. If preventative methods are appropriate, combining them with screening techniques such as CODERS may further enhance quality and reliability of the collected data for subsequent analysis.
Discussion
We introduced a novel method, CODERS, to identify the onset of careless responding (or absence thereof) for each participant in a given survey. It is motivated by literature stressing that in lengthy surveys—which are common in the behavioral and organizational sciences—a large proportion of survey participants may respond partially carelessly due to fatigue or boredom. CODERS is highly flexible and can combine multiple indicators for the potential onset of carelessness, such as autoencoder reconstruction errors as an indicator of inconsistency, and the proposed LongstringPattern (LSP) sequence as an indicator of invariability. Re-analyzing data from a seminal study on careless responding, we find evidence of partial careless responding that was hitherto undetected. Additionally, simulation experiments indicate excellent performance of CODERS in identifying the onset of careless responding, and highlight the importance of incorporating multiple indicators that capture different manifestations of carelessness.
While CODERS is a promising first step towards detecting periods of careless responding rather than identifying which respondents are careless, various pertinent questions remain for future research. For instance, the LSP sequence as a measurement of response invariability may require additional validation, as does the incorporation of response times into CODERS, which we leave for empirical follow-up work. In addition, extending or adapting CODERS with changepoint detection methods that allow for multiple changepoints, or that retain higher power in shorter surveys, may be fruitful directions for further research.
Whereas we apply machine learning and nonparametric techniques in our approach, other literature is focused on model-based approaches. For instance, Ulitzsch et al., 2022b propose a mixture model based on item response theory that is designed to detect careless respondents, which is adapted in Ulitzsch et al., 2022a to incorporate response times. Other approaches that explicitly model carelessness as part of a mixture model are proposed in Arias et al., (2020), Steinmann et al., (2022), and Van Laar & Braeken, (2022). One advantage of model-based approaches is that their estimates can be used to construct person-weights that can be used to downweight the adverse influence of potentially careless respondents in subsequent analyses, which is discussed in detail in Ulitzsch et al., 2022a ; Ulitzsch et al., 2022b and Hong & Cheng, (2019). However, model-based approaches typically impose strong and potentially unverifiable assumptions on the properties of attentive and careless responding which might be violated in practice.
On a final note, Arthur et al., (2021) envisage “artificial intelligence and machine learning being used to detect careless responding and response distortion in real time in the foreseeable future.” Using machine learning techniques in the form of autoencoders and nonparametric changepoint detection, our proposed method CODERS takes a step towards that future.
Computational Details
Throughout this paper, we have followed recent guidelines by Alfons & Welz, (2024) on the application of machine learning methods for the detection of careless responding. All computations were performed with R version 4.3.3 (R Core Team,, 2024) and Python version 3.10 (Van Rossum & Drake,, 2009). Our implementation of CODERS in the R package carelessonset is available from https://github.com/mwelz/carelessonset. Replication files will be made publicly available upon acceptance of this manuscript and a link will be included here.
Acknowledgments
We thank Donald Bergh, Dennis Fok, Patrick Groenen, Nick Koning, Robin Lumsdaine, Jakob Raymakers, and Kevin ten Haaf for valuable comments and feedback. This work was supported by a grant from the Dutch Research Council (NWO), research program Vidi (Project No. VI.Vidi.195.141).
Appendix A Setup and Assumptions
Throughout our paper, we consider the following setup. Let $\bm{X}$ be an $n\times p$ data matrix holding the rating-scale responses of $n$ survey participants to $p$ items. We make the standard assumption that all observations (i.e., all participants whose responses the survey collects) are independently and identically distributed. Moreover, it is a priori unknown if and when careless responding occurs in $\bm{X}$. We require the following assumptions.
Assumption 1.
The responses in $\bm{X}$ admit an $n\times s$ lower-dimensional representation, $\bm{S}$, where $s\ll p$. The dimension $s$ is known and corresponds to the number of constructs that the survey measures.
Assumption 2.
The survey that generated $\bm{X}$ is reliable in the sense that if all participants responded accurately, $\bm{X}$ would accurately measure all constructs.
Assumption 3.
The items in the survey are in a randomized order.
Assumption 4.
All participants begin the survey as attentive respondents by providing accurate and truthful responses. As the survey progresses, some (possibly none or all) participants start responding carelessly and continue to do so for the remainder of the survey.
Assumption 5.
The onset of partial carelessness is characterized by a changepoint in one or more indicators. We use autoencoder reconstruction errors and LongstringPattern (LSP) sequences as indicators.
Assumption 1 is very mild in survey data because surveys typically measure multiple constructs and the number of constructs is typically known. For instance, the $p=300$ item representation of the NEO-PI-R inventory (Goldberg,, 1999) measures $s=30$ constructs.
Assumption 2 is again mild in survey data. There exist well-established and reliable survey measures for a large variety of variables. For instance, the International Personality Item Pool (IPIP; Goldberg,, 1999) is a pool of more than 250 personality scales. It is in general recommended to use well-established measures to ensure a highly reliable measurement (Babbie,, 2020, p. 150). The reliability of a measurement can be estimated in multiple ways, such as Cronbach’s alpha, the omega coefficient and variations thereof, as well as the $H$ coefficient (see, e.g., McNeish,, 2018, for a description of these and other reliability measures). Hence, Assumption 2 is satisfied when the survey data were collected with measures of high reliability.
Assumption 3 is common practice in surveys in the social and behavioral sciences. The randomized order of items prevents that recurring response patters emerge as a result of the survey design rather than due to invariable careless responding.
Assumption 4 is discussed in the section on limitations, and Assumption 5 is motivated in detail in the section introducing the proposed methodology.
Appendix B Technical Details on CODERS
B.1 Details on Autoencoders
B.1.1 Network Architecture and Estimation
Denote by $x_{ij}$ the response of the $i$-th participant to the $j$-th survey item. Collect the responses of the $i$-th participant to all $p$ items in a $p$-dimensional vector $\mathbf{x}_{i}=(x_{i1},x_{i2},\dots,x_{ip})^{\top}$so that the $n\times p$ data matrix $\bm{X}$ is given by $\bm{X}=(\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{n})^{\top}$.
Introduced by Kramer, (1992), an autoencoder is a network consisting of an uneven number of $M$ layers, in which the first layer (input layer) holds the input data $\mathbf{x}_{i}$ and the last layer (output layer) holds the reconstructed input data $\widehat{\mathbf{x}}_{i}$. Denoting the number of nodes in the $\ell$-th layer by $N^{(\ell)}$, we have that $N^{(1)}=N^{(M)}=p$. Moreover, the central layer (bottleneck layer) consists of $s<p$ nodes to prevent that the input is simply passed through the network, and the network architecture is symmetric around the bottleneck layer.
Each node in a given layer stores so-called activations, which are obtained by transforming a linear combination of the activations from the previous layer. This transformation is defined by an activation function $g_{\ell}(\cdot)$. Formally, for participant $i=1,\dots,n$, the activation $a_{ij}^{(\ell)}$ of the $j$-th node, $j=1,\dots,N^{(\ell)},$ of the $\ell$-th layer, $\ell=2,\dots,M$, is given by
$a_{ij}^{(\ell)}$ $=g_{\ell}\left(z_{ij}^{(\ell)}\right),$
$z_{ij}^{(\ell)}$ $=\sum_{k=1}^{N^{(\ell-1)}}\omega_{jk}^{(\ell)}a_{ik}^{(\ell-1)}+b_{j}^{(\ell)}.$
The $\omega_{jk}^{(\ell)}$ and $b_{j}^{(\ell)}$ are unknown weights and intercept terms, respectively. Observe that in the first layer, we have $a_{ij}^{(1)}=x_{ij}$. The activations of the last layer, $a_{ij}^{(M)}$, hold the network’s output, namely the autoencoder’s reconstructions $\widehat{x}_{ij}$ of responses $x_{ij}$.
We jointly refer to the weights and intercept terms as the network’s parameters, which need to be estimated, and we collect them in a vector $\bm{\theta}$. To emphasize the dependence on the parameter vector $\bm{\theta}$, we define for a fixed $\bm{\theta}$ a prediction function $\bm{f}_{\bm{\theta}}(\cdot)$ of the network, which corresponds to the activations in the last layer:
$\bm{f}_{\bm{\theta}}(\mathbf{x}_{i})=\Big{(}f_{1,\bm{\theta}}(\mathbf{x}_{i}),f_{2,\bm{\theta}}(\mathbf{x}_{i}),\dots,f_{p,\bm{\theta}}(\mathbf{x}_{i})\Big{)}^{\top}=\left(a_{i1}^{(M)},a_{i2}^{(M)},\dots,a_{ip}^{(M)}\right)^{\top},$
for participants $i=1,\dots,n$. The reconstruction of the input vector $\mathbf{x}_{i}$ is then given by $\widehat{\mathbf{x}}_{i}~{}=~{}\bm{f}_{\bm{\theta}}(\mathbf{x}_{i})$. Due to the fact that each node is a function of nodes from a previous layer, the nodes in adjacent layers are often visualized by connecting edges. Figure B.1 provides such a schematic overview of a general autoencoder neural network.
Figure B.1: Schematic overview of an autoencoder neural network with $M$ layers.
[…]
Note. The first and last layer are referred to as input and output layer, respectively. For participants $i=1,\dots,n$, the input in the first layer are the input responses $\mathbf{x}_{i}=(x_{i1},\dots,x_{ip})^{\top}$, and the output in the last layer are the activations $f_{j,\bm{\theta}}(\mathbf{x}_{i})=a_{ij}^{(M)},\ j=1,\dots,p,$ that reconstruct the input responses for a parameter vector $\bm{\theta}$ to be estimated. The $j$-th node in the $\ell$-th layer holds the activation $a_{ij}^{(\ell)}$, for $\ell=2,\dots,M,j=1,\dots,N^{(\ell)}$. Besides the parameter vector $\bm{\theta}$, the prediction function $\bm{f}_{\bm{\theta}}(\cdot)$ depends implicitly on the prespecified choice of the activation functions $g_{\ell}(\cdot)$.
To fit an autoencoder, we aim at minimizing its reconstruction error. For a prespecified loss function $\mathcal{L}(\cdot)$, we fit the neural network by finding the $\bm{\theta}$ that yields the best average reconstruction error,
$\widehat{\bm{\theta}}=\arg\min_{\bm{\theta}}\Bigg{\{}\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{p}\mathcal{L}\Big{(}x_{ij}-f_{j,\bm{\theta}}(\mathbf{x}_{i})\Big{)}\Bigg{\}}.$ (B.1)
The prediction function of the fitted network is subsequently given by $\bm{f}_{\widehat{\bm{\theta}}}(\cdot)$. For the loss function in (B.1), we choose the smooth and robust Pseudo-Huber loss, which is defined for a fixed $\delta>0$ as
$\mathcal{L}(z)=\delta^{2}\left(\sqrt{1+(z/\delta)^{2}}-1\right).$ (B.2)
This choice results in quadratic loss for small values of $z$ and linear loss for large values of $z$. Consequently, the Pseudo-Huber loss function avoids that large individual reconstruction errors strongly affect the fit, which would make it hard for the network to distinguish attentive responses from inconsistent careless responses, as the latter typically lead to large prediction errors. To solve the optimization problem in (B.1), we use stochastic gradient descent. We refer to Chapter 8.5 in Goodfellow et al., (2016) for details.
B.1.2 Modeling Choices
Following Kramer, (1992), we specify $M=5$ layers in the autoencoder. We set the number of nodes in layers 2 and 4 (mapping an demapping layer, respectively) to $\lfloor 1.5\times p\rfloor$, as a relatively large number of nodes is expected to give the autoencoder flexibility to learn many different types of response behavior (attentive and careless). Based on Assumption 1, the number of nodes in the bottleneck layer equals the number of constructs measured in the questionnaire.
Concerning the activation functions, we again follow a recommendation in Kramer, (1992) and propose to use nonlinear activation functions in the mapping and demapping layers, as well as a linear activation function in the bottleneck layer. Specifically, we propose to use the hyperbolic tangent activation in the (de-)mapping layers and the identity mapping in the bottleneck layer. Definitions are given in Table B.1, which summarizes our proposed autoencoder architecture.
Table B.1: Choices regarding the network architecture of the autoencoder.
Layer Number of nodes Activation function
Mapping layer $\lfloor 1.5\times p\rfloor$ $\textsf{tanh}(x)=2/(1+\exp(-2x))-1$
Bottleneck layer $s$ $\textsf{identity}(x)=x$
Demapping layer $\lfloor 1.5\times p\rfloor$ $\textsf{tanh}(x)=2/(1+\exp(-2x))-1$
Note. The constants $p$ and $s$ denote the number of items in the questionnaire and the number of constructs that the questionnaire measures, respectively. For a real number $x$, $\lfloor x\rfloor$ denotes the largest integer smaller than or equal to $x$.
For fitting the autoencoder, we use the stochastic gradient descent algorithm with a batch size of 10, a learning rate of 0.0001, and 100 epochs. In the presudo-Huber loss (B.2), we set constant $\delta=1$.
B.2 LongstringPattern Algorithm
Formal algorithms for computing the $l$-pattern and LSP sequences are provided in Algorithms 1 and 2, respectively.
Input: Series of observed rating-scale responses $\{x_{j}\}_{j=1}^{p}$, pattern length $l\in\mathbb{N}$ to consider (number of items that an individual pattern encompasses). For instance, the individual pattern $1-2-3$ encompasses $l=3$ items in the response sequence $-3-1-2-$ 3 - 1 - 2 - 3. Note that the algorithm measures straightlining when setting $l=1$.
Output: Sequence of measures $\{l\textsf{-pattern}_{j}\}_{j=1}^{p}$
Initialization: $streak\leftarrow 1,startindex\leftarrow 1$
1
/* fill in the first $l-1$ indices */
2 for $k\leftarrow 1$ to $p-l-1$ by $1$ do
3       if $x[k]==x[k+l]$ then
            increment $streak$ by 1 ;
             // streak is ongoing
4
5      else
             Assign value $streak$ to all elements in $\{l\textsf{-pattern}_{j}\}_{j=startindex}^{k}$ ;
             // streak broke
6             $startindex\leftarrow k+1$;
7             streak $\leftarrow 1$;
8
9       end if
10
11 end for
12
/* fill in the last $l$ indices (there are 3 edge cases) */
13 if $\texttt{streak}\geq l$ then
       /* case 1: current streak is ongoing and we have seen it completely at least once, so assume it will last until the end and complete the current streak (which has not yet been broken) */
14       Assign value $streak+l-1$ to all elements in $\{l\textsf{-pattern}_{j}\}_{j=startindex}^{p}$;
15
16 else if $1<\texttt{streak}<l$ then
       /* case 2: there is an ongoing streak, but we have no chance of seeing it completely (since streak < l), so we do not know how the streak looks like. In this case, fill in the current value of "streak" until the (p-l)-th item and fill in 1 for the last l items (because we do not know what the complete streak looks like) */
17       Assign value $streak-1$ to all elements in $\{l\textsf{-pattern}_{j}\}_{j=startindex}^{p-l}$;
18       Assign value $1$ to all elements in $\{l\textsf{-pattern}_{j}\}_{j=p-l+1}^{p}$;
19
20 else
       /* case 3: there is no ongoing streak, so fill in 1 for last $l$ indices */
21       Assign value 1 to all elements in $\{l\textsf{-pattern}_{j}\}_{j=p-l+1}^{p}$;
22
23 end if
Algorithm 1 $l$-pattern($\{x_{j}\}_{j=1}^{p},l$)
Input: Series of observed rating-scale responses $\{x_{j}\}_{j=1}^{p}$, maximum pattern length $l_{max}\in\mathbb{N}$ to consider
Output: Set of LongstringPattern measures $\{\textsf{LSP}_{j}\}_{j=1}^{p}$
1
/* calculate $l$-pattern for fixed pattern length $l$/
2 for $l\leftarrow 1$ to $l_{max}$ by $1$ do
3       $\{\textsf{LSP}^{l}_{j}\}_{j=1}^{p}\leftarrow l\textsf{-pattern}(\{x_{j}\}_{j=1}^{p},l)$;
4
5 end for
6
/* for each element $\{1,\dots,p\}$, report the largest $l$-pattern value */
7 Initialize $\{\textsf{LSP}_{j}\}_{j=1}^{p}$;
8 for $j\leftarrow 1$ to $p$ by $1$ do
9       $\textsf{LSP}_{j}\leftarrow\max\big{\{}\textsf{LSP}_{j}^{1},\textsf{LSP}_{j}^{2},\dots,\textsf{LSP}_{j}^{l}\big{\}}$;
10
11 end for
Algorithm 2 LongstringPattern($\{x_{j}\}_{j=1}^{p},l_{max}$)
B.3 Details on Changepoint Detection
Let $\{\bm{Y}_{j}\}_{j=1}^{p}$ be a series of length $p$ consisting of $d$-dimensional random variables. The goal of the cumulative sum self-normalization test of Shao & Zhang, (2010) is to estimate the location of a possible changepoint in the (multivariate) series, i.e., to divide the series into two subsets $\{\bm{Y}_{j}\}_{j=1}^{k-1}$ and $\{\bm{Y}_{j}\}_{j=k}^{p}$ in which a $d$-dimensional parameter of interest $\bm{\theta}$ differs. Here, we focus on the mean as the parameter $\bm{\theta}$. Define by $\widehat{\bm{\theta}}_{a,b}=(b-a+1)^{-1}\sum_{j=a}^{b}\bm{Y}_{j}$ the mean of the series calculated on the subset implied by periods $a\leq b$. For location $k\in\{2,\dots,p\}$, define the statistic
$T_{p}(k)=\bm{D}_{p}(k)^{\top}\bm{V}_{p}(k)^{-1}\bm{D}_{p}(k),$
where
$\bm{D}_{p}(k)$ $=\frac{(k-1)(p-k+1)}{p^{3/2}}\left(\widehat{\bm{\theta}}_{1,k-1}-\widehat{\bm{\theta}}_{k,p}\right),$
$\bm{V}_{p}(k)$ $=\bm{L}_{p}(k)+\bm{R}_{p}(k),$
with
$\bm{L}_{p}(k)=\left\{\begin{array}[]{ll}\displaystyle\sum_{i=1}^{k-2}\frac{i^{2}(k-1-i)^{2}}{p^{2}(k-1)^{2}}\left(\widehat{\bm{\theta}}_{1,i}-\widehat{\bm{\theta}}_{i+1,k-1}\right)\left(\widehat{\bm{\theta}}_{1,i}-\widehat{\bm{\theta}}_{i+1,k-1}\right)^{\top},&\qquad\text{if }k>2,\\ 0,&\qquad\text{if }k=2,\\ \end{array}\right.$
and
$\bm{R}_{p}(k)=\left\{\begin{array}[]{ll}\displaystyle\sum_{i=k+1}^{p}\frac{(p-i+1)^{2}(i-k)^{2}}{p^{2}(k-1)^{2}}\left(\widehat{\bm{\theta}}_{i,p}-\widehat{\bm{\theta}}_{k,i-1}\right)\left(\widehat{\bm{\theta}}_{i,p}-\widehat{\bm{\theta}}_{k,i-1}\right)^{\top},&\qquad\text{if }k<p,\\ 0,&\qquad\text{if }k=p,\\ \end{array}\right.$
The final test statistic is then defined as
$SN_{p}=\max_{k=2,\dots,p}T_{p}(k).$ (B.3)
Consider a given threshold $K_{d}>0$. If the test statistic exceeds the threshold, $SN_{p}>K_{d}$, an estimate of the changepoint location is obtained via
$\widehat{k}=\operatorname*{arg\,max}_{k=2,\dots,p}T_{p}(k).$
Otherwise, no changepoint is detected. Shao & Zhang, (2010) and Zhao et al., (2022) derive theoretical guarantees for this procedure. Specifically, Shao & Zhang, (2010) obtain the asymptotic distribution of the test statistic $SN_{p}$ for $p\rightarrow\infty$ and provide the corresponding critical values $K_{d}$ for a given significance level $\alpha$ and dimension $d$. We summarize relevant critical values in Table B.2.
Table B.2: Critical values $K_{d}$ for relevant significance levels $\alpha$ and dimensions $d$.
Dimension $d$
$\alpha$ 1 2
0.010 68.6 117.7
0.005 84.6 135.3
0.001 121.9 192.5
Note. The critical values are taken from Table 1 in Shao & Zhang, (2010).
In the context of this paper, we use the test statistic from (B.3) to test for changepoints in the two-dimensional series
$\bm{Y}_{j}=\begin{pmatrix}\textsf{RE}_{j}\\ \textsf{LSP}_{j}\end{pmatrix},\qquad j=1,\dots,p,$
where $\textsf{RE}_{j}$ is the autoencoder reconstruction error from (1) of a response to item $j$, and $\textsf{LSP}_{j}$ is the value of the LSP sequence from (2) assigned to this response. If a changepoint $\widehat{k}$ is detected, under Assumptions 4 and 5, the subsets $\{\bm{Y}_{j}\}_{j=1}^{\widehat{k}-1}$ and $\{\bm{Y}_{j}\}_{j=\widehat{k}}^{p}$ may be interpreted as attentive and careless responses, respectively. That is, the detected changepoint $\widehat{k}$ may be viewed as the onset of careless responding. We obtain such a series for each of the $n$ survey participants and subsequently test each of the $n$ series for changepoints.
On a final technical note, LSP sequences may exhibit low variation. Computing their variation over a relatively narrow interval—as required by the test statistic $SN_{p}$—may result in a singular (and therefore not invertible) variation matrix $\bm{V}_{p}(k)$ so that $SN_{p}$ cannot be computed. To avoid this issue, we by default inject Gaussian noise of tiny magnitude to each $\textsf{LSP}_{j},j=1,\dots,p$. Specifically, we inject draws from a normal distribution with mean zero and variance $0.01^{2}$. Doing so results in negligible but nonzero variation, which renders $\bm{V}_{p}(k)$ invertible.
Appendix C Details on the Empirical Application
C.1 Data Preprocessing
The data of Johnson, (2005), provided at https://osf.io/sxeq5/, contain the following information for each of the the $n=20,993$ respondents. First, per-item responses to $p=300$ IPIP-NEO-300 items, five domain scores (directly calculated from the responses), scores for psychometric antonyms and personal reliability, as well as respondent sex and age.
The responses in the provided data have already underwent some preprocessing that is described in detail the accompanying README file. Specifically, responses to negatively-worded items were reverse-coded, that is, a given response of 5 was recoded to 1, 4 was recoded to 3, and so on. Hence, in order to recover the original responses, we reverted this recoding. Furthermore, a total of 33,691 individual responses are missing (0.53% of all responses). Note that Johnson, had already removed participants with too many missing responses from the provided data set, hence there are relatively few missing responses for the remaining participants. Due to the low number of missing responses, we treat them as a separate answer category by coding them as 0.
C.2 Additional Results
Figure C.1 visualizes the detected onset items for the respondents that are flagged as careless by CODERS at significance levels 0.1%, 0.5%, and 1%.
Figure C.1: Estimated onset of careless responding for participants flagged by CODERS in the data of Johnson, (2005).
Note. Results for different significance levels $\alpha\in\{0.01,0.05,0.1\}$ are presented in separate panels. The $x$-axis shows the indices of the participants flagged as partially carelessness by CODERS. The participants are sorted in ascending order of the estimated carelessness onset item, which is in turn depicted on the $y$-axis. The dashed horizontal line represents the average location of carelessness onset across the flagged participants.
Figure C.2 provides a histogram as well as a density estimate for the per-respondent average reconstruction errors. We can see that the density has a long right tail, which might be due to careless respondents. We highlight the value 0.064 by a vertical dotted line, which is the average reconstruction error of the participant from Figure 3(d) and Table 2(d). This value is in the 99.7% percentile of all $n=20,993$ average reconstruction errors. Given the extremely high average reconstruction error and the fact that this respondent was flagged by both benchmark methods, it is likely that the participant responded carelessly throughout the survey, explaining why CODERS did not flag a changepoint.
Figure C.2: Histogram of average reconstruction errors for the data of Johnson, (2005).
Note. The $x$-axis contains the per-respondent average reconstruction errors, $\textsf{RE}_{i}=\frac{1}{p}\sum_{j=1}^{p}\textsf{RE}_{ij}$, $i=1,\dots,n$, where averages are taken over the $p=300$ items. On the $y$-axis, a histogram of the average reconstruction errors for $n=20,993$ respondents is displayed together with an empirical Gaussian kernel density. The vertical dotted line denotes the value of $\textsf{RE}_{i}$ for the participant in Figure 3(d) that was flagged by both benchmark methods, but not CODERS. This value, 0.064, lies in the 99.7% percentile of all average reconstruction errors.
Appendix D Details and Additional Simulations
D.1 Details on the Adjusted Rand Index
In the context of our simulations, the adjusted Rand index (ARI) for a given respondent is computed as
$ARI=\begin{cases}1&\text{ if the respondent is attentive and no changepoint isflagged},\\ \frac{A-BC/D}{(B+C)/2-BC/D}&\textnormal{ otherwise,}\end{cases}$
where
$A$ $=\binom{p_{11}}{2}+\binom{p_{00}}{2}+\binom{p_{10}}{2}+\binom{p_{01}}{2},$
$B$ $=\binom{p_{00}+p_{01}}{2}+\binom{p_{10}+p_{11}}{2},$
$C$ $=\binom{p_{00}+p_{10}}{2}+\binom{p_{01}+p_{11}}{2},$
$D$ $=\binom{p}{2},$
with $p_{11}$ denoting the number of careless responses that are correctly identified as careless, $p_{00}$ the number of attentive responses that are correctly identified as attentive, $p_{10}$ the number of careless responses that are incorrectly identified as attentive, and $p_{01}$ the number of attentive responses that are incorrectly identified as careless. Intuitively, $p_{11}$ and $p_{00}$ measure correct classification, whereas $p_{01}$ and $p_{10}$ measure misclassification. Observe that the ARI assumes value 1 if there is no misclassification, which is the maximum value it can take. Hence, for accurate estimates of the carelessness onset item, the ARI will be close to value 1. For very inaccurate estimates of carelessness onset, the ARI will be close to 0, or potentially even below that. It is unclear how to interpret negative ARI values, but such values only occur rarely. We refer to Steinley, (2004) for a detailed discussion of the properties of the ARI.
If a changepoint is flagged in an attentive respondent, the ARI takes value 0. Similarly, if no changepoint is flagged in an attentive respondent, the ARI takes value 1. Hence, when averaging the ARI over attentive respondents, the ensuing average can be interpreted as the proportion of attentive respondents who are correctly identified as attentive. On the other hand, averages of ARIs calculated on partially careless respondents do not have a direct interpretation other than higher values indicating better performance.
D.2 Additional Simulation Results
We investigate the effect of various choices in our simulations. As the results are similar to those in the main paper other than some shifts in size or power, we only present plots of the average ARI for attentive and careless respondents (similar to Figure 5) in order to keep the appendix concise.
D.2.1 Choosing a Different Significance Level
Figure D.1: ARI of CODERS at a significance level of 0.5%.
Note. The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows).
Figure D.2: ARI of CODERS at a significance level of 1%.
Note. The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows).
We use the baseline simulation design from the main text, but we set the significance level in CODERS to 0.5% and to 1%, respectively. The corresponding results are shown in Figures D.1 and D.2. Compared to the significance level of 0.1% in the main text (cf. Figure 5), there is a gradual increase in power in the changepoint detection, yielding a small increase in the ARI across prevalence levels for careless respondents. Further investigation suggests that the increase in ARI stems mainly from better performance in detecting random and extreme responding (results not shown to keep the appendix at a reasonable length). However, for high prevalence levels from 40% onwards, this comes at a cost of too many false positives in attentive respondents, translating to a substantial decrease in ARI for the highest prevalence levels.
D.2.2 Varying Correlation Structures
In this variation of the simulation design, we keep the same settings as in the baseline design from the main text, but we vary the correlation structure in the following ways:
  1. 1.
    Randomly sampled moderate correlations. In every repetition, correlations of item pairs from the same scale are randomly drawn from the interval $[0.4,0.6]$. Items from different scales remain uncorrelated.
  2. 2.
    Randomly sampled strong correlations. In every repetition, correlations of item pairs from the same scale are randomly drawn from the interval $[0.6,0.8]$. Items from different scales remain uncorrelated.
  3. 3.
    Randomly sampled grouped correlation structure. Multiple scales form groups so that there are 5 groups consisting of 6 scales each. Since there are 10 items in each scale, there are $6\times 10=60$ items in each group. In each repetition, correlations of item pairs from the same scale are randomly drawn from the interval $[0.6,0.8]$, correlations of item pairs in different scales of the same group are randomly drawn from the interval $[0,0.2]$, and items from different groups are uncorrelated.
While the baseline simulation design uses a relatively simple correlation structure (items from the same scale have a fixed correlation of 0.7 before recoding, items from different scales are uncorrelated), the three variations above introduce more randomness into the design. Due to this randomness, it may occur that a sampled correlation matrix is not positive definite. In such a case, we find the closest positive definite correlation matrix via the function nearPD() in the R package Matrix (Bates et al.,, 2023).
Figure D.3: ARI of CODERS for randomly sampled moderate correlations.
Note. The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows).
Figure D.4: ARI of CODERS for randomly sampled strong correlations.
Note. The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows).
Figure D.5: ARI of CODERS for the randomly sampled grouped correlation structure.
Note. The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows).
Figures D.3D.5 display the results corresponding to the different correlation structures at significance level 0.1%. The results of the design with randomly sampled strong correlations (Figure D.4) and the grouped correlation structure (Figure D.5) are essentially equivalent to the results of the baseline design with fixed correlation (Figure 5 in the main text). The design with randomly sampled moderate correlations (Figure D.3) also yields similar results, save for a small power drop in the changepoint detection (resulting in a lower ARI for careless respondents) and fewer false positives in extreme carelessness prevalence levels (yielding a higher ARI in attentive respondents).
D.2.3 Varying Sample Size
This variation of the simulation design again uses the same settings as in the baseline design from the main text, but now we change the sample size to $n=250$ and $n=1000$, respectively.
Figures D.6 and D.7 visualize the corresponding results at significance level 0.1%. They are similar to the results for $n=500$ in the main text (cf. Figure 5), other than a small decrease ARI in careless respondents for $n=250$ as well as a small increase for $n=1000$.
Figure D.6: ARI of CODERS for sample size $n=250$.
Note. The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows).
Figure D.7: ARI of CODERS for sample size $n=1000$.
Note. The ARI is averaged over fully attentive respondents and partially careless respondents (columns) for different carelessness onset settings (rows).